# Alternating Direction Method of Multipliers AR AutoRally

### v3

@@ II-C MPPI-Generic’s MPPI Implementation
// Begin Cost sampling
@@ 15 UnicycleCost(cudaStream_t stream = nullptr);
18 { return ”Unicycle Cost”; }
@@ V Benchmarks
The next implementation we compare against is ROS2’s MPPI. As of ROS Iron, there is a CPU implementation of MPPI in the ROS navigation stack [58]. This CPU implementation is written in C++ and looks to make heavy use of AVX or vectorized instructions to improve performance. There is a small selection of dynamics models (Differential Drive, Ackermann, and Omni-directional) and cost functions that are focused around wheeled robots navigating through obstacle-laden environments. This implementation will only become more widespread as ROS2 adoption continues to grow over the coming years, making it an essential benchmark. Unfortunately, it does have some drawbacks as it is not possible to add new dynamics or cost functions without rewriting the base code itself, has no implementation of Tube-MPPI or RMPPI, and is only available in ROS2 Iron or newer. This means that it might not be usable on existing hardware platforms that are unable to upgrade their systems.
@@
In order to compare our library against these three implementations, we recreated the same dynamics and cost function for each version of MPPI. As ROS2’s implementation would be the hardest to modify, we chose the Differential Drive dynamics model and some of the cost function components that already exist there as the baseline. We used the goal position quadratic cost, goal angle quadratic cost, and the costmap-based obstacle cost components so that we could maintain a fairly simple cost function that allows us to show the capabilities of our library. We implemented these dynamics and cost functions in both CUDA and Python. The CUDA implementations were extensions of our base Dynamics and Cost Function APIs. We decided to use the same code in the Autorally implementation as well which required some minor rewriting to account for different method names and state dimensions. The Python implementation was an extension of the TorchRL base Environment class, and used PyTorch’s JIT compiler to speed up performance when used in the TorchRL implementation. We used the same parameters for sampling, dynamics, cost function tuning, and MPPI hyperparameters across all implementations, summarized in Table I.
@@
As a result, the optimization times shown are not going to be the fastest possible performance that can be achieved on any given GPU but these tests should still serve as a useful benchmark to understand the average performance that can be achieved. Our test was timing how long each implementation of MPPI would take to return an optimal control sequence when provided an initial state, ${\bf x}_{0}$. We ran each of the Autorally, MPPI-Generic, and ROS2 implementations $10,000$ times to produce optimal trajectories with 128, 256, 512, 1024, 2048, 4096, 6144, 8192, and 16,384 samples; the TorchRL implementation was only run $1000$ times due to it being too slow to compute, even when using the GPU. The comparisons were run across a variety of hardware including a Jetson Nano to see what bottlenecks each implementation might have. The Jetson Nano was unfortunately only able to run the MPPI-Generic and Autorally MPPI implementations as the last supported PyTorch version and the lastest TorchRL libraries were incompatible, and the ROS2 implementation was unable to compile. GPUs tested ranged from a NVIDIA GTX 1050 Ti to a NVIDIA RTX 4090. Most tests were performed on an Intel 13900K which is one of the fastest available CPUs at the time of this writing in order to prevent the CPU being the bottleneck for the mostly GPU-based comparison; however, we also ran tests on an AMD Ryzen 5 5600x to see the difference in performance on a lower-end CPU. The MPPI optimization times across all hardware can be seen in Table III. The code used to do these comparisons is available at https://github.com/ACDSLab/MPPI_Paper_Example_Code.
@@ V-A Results
Figure 3: Optimization times for all MPPI implementations on a hardware system with a RTX 4090 and an Intel 13900K over a variety of number of samples. The ROS2 CPU implementation grows linearly as the number of samples increase while GPU implementations grow more slowly.
Going over all of the collected data would take too much room for this paper so we shall instead try to pull out interesting highlights to discuss. Full results can be seen in Table III. First, we look at the results on the most powerful system tested, using an Intel i9-13900K and an NVIDIA RTX 4090 in Fig. 3. As the number of samples increase, the CPU-bound ROS2 method increases in optimization times in a linear fashion. Every other method uses the GPU and we see little reason to use small sample sizes as they have the same computation time till we reach around 1024 samples. As we hit $16,384$ samples, the AutoRally implementation starts to have lower optimization times than MPPI-Generic. We will see this trend continue in Fig. 4.
Figure 4: Optimization times for all MPPI implementations on a hardware system with a GTX 1050 Ti and an Intel 13900K over a variety of number of samples. MPPI-Generic and AutoRally on this older hardware eventually start to scale linearly with the number of samples but does so at a much lower rate with our library compared to ROS2 or TorchRL.
When looking at older and lower-end NVIDIA hardware such as the GTX 1050 Ti, our library still performs well compared to other implementations as seen in Fig. 4. Only when the number of samples is at 128 does the ROS2 implementation on an Intel 13900k match the performance of the AutoRally implementation on this older GPU. MPPI-Generic is still more performant at these lower number of samples and eventually it scales linearly as we get to thousands of samples. The TorchRL implementation also finally starts to show some GPU bottle-necking as we start to see optimization times increasing as we reach over 6144 samples. There is also a moment where the MPPI-Generic library optimization time grows to be larger than the AutoRally implementation. That occurs when we switch from using the split kernels (Section III-C2) to the combined kernel (Section III-C3). The AutoRally implementation uses a combined kernel with fewer GPU synchronization points due to strictly requiring forward Euler integration for the dynamics. At the small hit to performance in the combined kernel, our library allows for many more features, such as multi-threaded cost functions, use of shared memory in the cost function, and implementation of more computationally-heavy integration methods such as Runge-Kutta or backward Euler integration. And while we see a hit to performance when using the combined kernel compared to AutoRally, we still see that the split kernel is faster for up to 2048 samples.
@@
We also conducted tests on a Jetson Nano to show that even on relatively low-power and older systems, our library can still be used. As the latest version of CUDA supporting the Jetson Nano is 10.2 and the OS is Ubuntu 18.04, both the TorchRL and ROS2 MPPI implementations were not compatible. As such, we only have results comparing our MPPI-Generic implementation to the AutoRally implementation in Fig. 6. Here, the AutoRally implementation starts having faster compute times around 512 samples. Again, this is due to our library switching to the combined kernel which will be slower. However, our library on a Jetson Nano at 2048 samples has a roughly equivalent computation time to that of 2048 samples of the ROS2 implementation on the Intel 13900K process, showing that our GPU parallelization can allow for real-time optimization even on portable systems.
@@
In addition, we see the benefits of our library as we increase the computation time of the cost function. At this point, the TorchRL and ROS2 implementations have been shown to be slow in comparison to the other implementations and are thus dropped from this cost complexity comparison.
@@ VII Acknowledgements
AMD 5600X 8192 ros2 $40.6365\pm 1.059\,23$ Intel 13900K 3080 8192 torchrl $28.978\,863\pm 1.443\,538\,958$
AMD 5600X 3080 8192 torchrl $54.198\,904\,75\pm 2.093\,610\,932$ Intel 13900K 8192 ros2 $29.9169\pm 0.239\,302$
@@
Jetson Orin Nano (8 GB) 128 autorally $0.946\,763\pm 0.292\,363$ Jetson Orin Nano (8 GB) 128 MPPI-Generic $0.629\,243\pm 0.1296$
Jetson Orin Nano (8 GB) 256 autorally $0.903\,75\pm 0.066\,688\,6$ Jetson Orin Nano (8 GB) 256 MPPI-Generic $0.607\,396\pm 0.022\,103\,4$
Jetson Orin Nano (8 GB) 512 autorally 0.909 $09\,514\pm 0.033\,216\,3$ Jetson Orin Nano (8 GB) 512 MPPI-Generic $0.728\,221\pm 0.030\,154\,5$
Jetson Orin Nano (8 GB) 1024 autorally $1.031\,62\pm 0.030\,783\,8$ Jetson Orin Nano (8 GB) 1024 MPPI-Generic $1.030\,02\pm 0.034\,969\,4$
Jetson Orin Nano (8 GB) 2048 autorally $1.408\,94\pm 0.055\,614\,6$ Jetson Orin Nano (8 GB) 2048 MPPI-Generic $1.645\,41\pm 0.108\,883$
Jetson Orin Nano (8 GB) 4096 autorally $2.712\,73\pm 0.124\,459$ Jetson Orin Nano (8 GB) 4096 MPPI-Generic $3.092\,44\pm 0.079\,106\,5$
Jetson Orin Nano (8 GB) 6144 autorally $3.862\,98\pm 0.100\,405$ Jetson Orin Nano (8 GB) 6144 MPPI-Generic 4.545 $45\,14\pm 0.135\,722$
Jetson Orin Nano (8 GB) 8192 autorally $5.255\,01\pm 0.131\,005$ Jetson Orin Nano (8 GB) 8192 MPPI-Generic $6.102\,01\pm 0.153\,504$
Jetson Orin Nano (8 GB) 16384 autorally $9.940\,03\pm 0.196\,983$ Jetson Orin Nano (8 GB) 16384 MPPI-Generic $12.0695\pm 0.217\,526$

### v4

@@ II-C MPPI-Generic’s MPPI Implementation
1
1ex// Begin Cost sampling
@@ 15 UnicycleCost(cudaStream_t stream = nullptr);
18 { return "Unicycle Cost"; }
@@ V Benchmarks
The next implementation we compare against is Nav2’s MPPI. As of ROS Humble, there is a CPU implementation of MPPI in the ROS navigation stack [30]; for our testing, we utilized the version packaged for ROS 2 Iron. This CPU implementation [7] is written in C++ and looks to make heavy use of AVX or other vectorized instructions to improve performance. There is a small selection of dynamics models (Differential Drive, Ackermann, and Omni-directional) and cost functions that are focused around wheeled robots navigating through obstacle-laden environments. It allows users to implement new cost functions as critics through a plugin system and new dynamics models based off its MotionModel base class. This implementation will only become more widespread as ROS2 adoption continues to grow over the coming years, making it an essential benchmark. Unfortunately, it does have some drawbacks as it has no implementation of Tube-MPPI or RMPPI, and is only available in ROS2 Humble or newer. This means that it might not be usable on existing hardware platforms that are unable to upgrade their systems.
@@
In order to compare our library against these three implementations, we recreated the same dynamics and cost function for each version of MPPI. We chose the Differential Drive dynamics model and some of the cost function components that already exist in the Nav2 impelemtation as the baseline. We used the goal position quadratic cost, goal angle quadratic cost, and the costmap-based obstacle cost components so that we could maintain a fairly simple cost function that allows us to show the capabilities of our library. We implemented these dynamics and cost functions in both CUDA and Python. The CUDA implementations were extensions of our base Dynamics and Cost Function APIs. We decided to use the same code in the Autorally implementation as well which required some minor rewriting to account for different method names and state dimensions. The Python implementation was an extension of the TorchRL base Environment class, and used PyTorch’s JIT compiler to speed up performance when used in the TorchRL implementation. We used the same parameters for sampling, dynamics, cost function tuning, and MPPI hyperparameters across all implementations, summarized in Table I.
@@
As a result, the optimization times shown are not going to be the fastest possible performance that can be achieved on any given GPU but these tests should still serve as a useful benchmark to understand the average performance that can be achieved. Our test was timing how long each implementation of MPPI would take to return an optimal control sequence when provided an initial state, ${\bf x}_{0}$. We ran each of the Autorally, MPPI-Generic, and Nav2 implementations $10,000$ times to produce optimal trajectories with 128, 256, 512, 1024, 2048, 4096, 6144, 8192, and 16,384 samples; the TorchRL implementation was only run $1000$ times due to it being too slow to compute, even when using the GPU. The comparisons were run across a variety of hardware including a Jetson Nano to see what bottlenecks each implementation might have. The Jetson Nano was unfortunately only able to run the MPPI-Generic and Autorally MPPI implementations as the last supported PyTorch version and the lastest TorchRL libraries were incompatible, and the Nav2 implementation was unable to compile. GPUs tested ranged from a NVIDIA GTX 1050 Ti to a NVIDIA RTX 4090. Most tests were performed on an Intel 13900K which is one of the fastest available CPUs at the time of this writing in order to prevent the CPU being the bottleneck for the mostly GPU-based comparison; however, we also ran tests on an AMD Ryzen 5 5600x to see the difference in performance on a lower-end CPU. The MPPI optimization times across all hardware can be seen in Table III. The code used to do these comparisons is available at https://github.com/ACDSLab/MPPI_Paper_Example_Code.
@@ V-A Results
Figure 3: Optimization times for all MPPI implementations on a hardware system with a RTX 4090 and an Intel 13900K over a variety of number of samples. The Nav2 CPU implementation grows linearly as the number of samples increase while GPU implementations grow more slowly.
@@
Going over all of the collected data would take too much room for this paper so we shall instead try to pull out interesting highlights to discuss. Full results can be seen in Table III. First, we look at the results on the most powerful system tested, using an Intel i9-13900K and an NVIDIA RTX 4090 in Fig. 3. As the number of samples increase, the CPU-bound Nav2 method increases in optimization times in a linear fashion immediately. Every other method uses the GPU which can provide headroom to increase the number of samples as the GPU is not fully utilized at small sample sizes. As we hit $16,384$ samples, the AutoRally implementation starts to have lower optimization times than MPPI-Generic. We will see this trend continue in Fig. 4.
Figure 4: Optimization times for all MPPI implementations on a hardware system with a GTX 1050 Ti and an Intel 13900K over a variety of number of samples. MPPI-Generic and AutoRally on this older hardware eventually start to scale linearly with the number of samples but does so at a much lower rate with our library compared to Nav2 or TorchRL.
When looking at older and lower-end NVIDIA hardware such as the GTX 1050 Ti, our library still performs well compared to other implementations as seen in Fig. 4. Only when the number of samples is at 128 does the Nav2 implementation on an Intel 13900k match the performance of the AutoRally implementation on this older GPU. However, we can see that Nav2 implementation again outperforms the TorchRL implementation until we hit 16,384 samples. MPPI-Generic is still more performant at these lower number of samples and eventually it scales linearly as we get to thousands of samples. The TorchRL implementation also finally starts to show some GPU bottle-necking as we start to see optimization times increasing as we reach over 6144 samples. There is also a moment where the MPPI-Generic library optimization time grows to be larger than the AutoRally implementation. That occurs when we switch from using the split kernels (Section III-C2) to the combined kernel (Section III-C3). The AutoRally implementation uses a combined kernel with fewer GPU synchronization points due to strictly requiring forward Euler integration for the dynamics. At the small hit to performance in the combined kernel, our library allows for many more features, such as multi-threaded cost functions, use of shared memory in the cost function, and implementation of more computationally-heavy integration methods such as Runge-Kutta or backward Euler integration. And while we see a hit to performance when using the combined kernel compared to AutoRally, we still see that the split kernel is faster for up to 2048 samples.
@@
We also conducted tests on a 2019 Jetson Nano to show that even on relatively low-power and older systems, our library can still be used. As the latest version of CUDA supporting the Jetson Nano is 10.2 and the OS is Ubuntu 18.04, both the TorchRL and Nav2 MPPI implementations were not compatible. As such, we only have results comparing our MPPI-Generic implementation to the AutoRally implementation in Fig. 6. Here, the AutoRally implementation starts having faster compute times around 512 samples. Again, this is due to our library switching to the combined kernel which will be slower. However, our library on a Jetson Nano at 2048 samples has a roughly equivalent computation time to that of 2048 samples of the Nav2 implementation on the Intel 13900K process, showing that our GPU parallelization can allow for real-time optimization even on portable systems.
@@
In addition, we see the benefits of our library as we increase the computation time of the cost function. At this point, the TorchRL and Nav2 implementations have been shown to be slow in comparison to the other implementations and are thus dropped from this cost complexity comparison.
@@ VII Acknowledgements
AMD 5600X 8192 nav2 $40.6365\pm 1.059\,23$ Intel 13900K 3080 8192 torchrl $28.978\,863\pm 1.443\,538\,958$
AMD 5600X 3080 8192 torchrl $54.198\,904\,75\pm 2.093\,610\,932$ Intel 13900K 8192 nav2 $29.9169\pm 0.239\,302$
@@
Jetson Orin Nano (8 GB) 128 MPPI-Generic $0.629\,243\pm 0.1296$ Jetson Orin Nano (8 GB) 4096 autorally $2.712\,73\pm 0.124\,459$
Jetson Orin Nano (8 GB) 128 autorally $0.946\,763\pm 0.292\,363$ Jetson Orin Nano (8 GB) 4096 MPPI-Generic $3.092\,44\pm 0.079\,106\,5$
Jetson Orin Nano (8 GB) 128 nav2 $2.707\,37\pm 0.021\,321\,2$ Jetson Orin Nano (8 GB) 4096 nav2 $87.3834\pm 0.260\,187$
Jetson Orin Nano (8 GB) 128 torchrl $244.523\,53\pm 6.836\,38$ Jetson Orin Nano (8 GB) 4096 torchrl $278.757\,93\pm 6.935\,90$
Jetson Orin Nano (8 GB) 256 MPPI-Generic $0.607\,396\pm 0.022\,103\,4$ Jetson Orin Nano (8 GB) 6144 autorally $3.862\,98\pm 0.100\,405$
Jetson Orin Nano (8 GB) 256 autorally $0.903\,75\pm 0.066\,688\,6$ Jetson Orin Nano (8 GB) 6144 MPPI-Generic 4.545 $45\,14\pm 0.135\,722$
Jetson Orin Nano (8 GB) 256 nav2 $5.148\,64\pm 0.063\,803\,6$ Jetson Orin Nano (8 GB) 6144 nav2 $129.927\pm 0.347\,753$
Jetson Orin Nano (8 GB) 256 torchrl $272.142\,35\pm 8.284\,30$ Jetson Orin Nano (8 GB) 6144 torchrl $283.596\,09\pm 8.088\,59$
Jetson Orin Nano (8 GB) 512 MPPI-Generic $0.728\,221\pm 0.030\,154\,5$ Jetson Orin Nano (8 GB) 8192 autorally $5.255\,01\pm 0.131\,005$
Jetson Orin Nano (8 GB) 512 autorally 0.909 $09\,514\pm 0.033\,216\,3$ Jetson Orin Nano (8 GB) 8192 MPPI-Generic $6.102\,01\pm 0.153\,504$
Jetson Orin Nano (8 GB) 512 nav2 $10.0513\pm 0.057\,195\,1$ Jetson Orin Nano (8 GB) 8192 nav2 $170.249\pm 0.356\,299$
Jetson Orin Nano (8 GB) 512 torchrl $272.298\,624\pm 6.714\,30$ Jetson Orin Nano (8 GB) 8192 torchrl $289.851\,03\pm 6.660\,31$
Jetson Orin Nano (8 GB) 1024 MPPI-Generic $1.030\,02\pm 0.034\,969\,4$ Jetson Orin Nano (8 GB) 16384 autorally $9.940\,03\pm 0.196\,983$
Jetson Orin Nano (8 GB) 1024 autorally $1.031\,62\pm 0.030\,783\,8$ Jetson Orin Nano (8 GB) 16384 MPPI-Generic $12.0695\pm 0.217\,526$
Jetson Orin Nano (8 GB) 1024 nav2 $273.906\,42\pm 8.264\,01$ Jetson Orin Nano (8 GB) 16384 torchrl $314.679\,59\pm 8.521\,67$
Jetson Orin Nano (8 GB) 1024 torchrl $21.4881\pm 0.105\,38$ Jetson Orin Nano (8 GB) 16384 nav2 $334.911\pm 0.745\,98$
Jetson Orin Nano (8 GB) 2048 autorally $1.408\,94\pm 0.055\,614\,6$
Jetson Orin Nano (8 GB) 2048 MPPI-Generic $1.645\,41\pm 0.108\,883$
Jetson Orin Nano (8 GB) 2048 nav2 $43.7964\pm 0.186\,287$
Jetson Orin Nano (8 GB) 2048 torchrl $275.457\,77\pm 8.236\,31$